OpenStudy (thedj4jc):

Solve (x + 1 < 5) ∩ (x - 4 > -3). {all real numbers} {x | 1 < x < 4} {x | x < 4 or x > 1}

OpenStudy (thedj4jc):

@BIGDADDY69420 @sstylesprettygirl @texaschic101 Will one of you see if you can help me on this?

OpenStudy (anonymous):

absolutely

OpenStudy (thedj4jc):

Thank you!! I have no idea how to solve this.

OpenStudy (anonymous):

@iGreen

OpenStudy (anonymous):

yes i will help u

OpenStudy (thedj4jc):

Yay, thanks!!

OpenStudy (anonymous):

so you want to solve for x right

OpenStudy (thedj4jc):

I think so

OpenStudy (anonymous):

so you want to add 1 to each side i think im not that good but heres what i got5x - 1 +1 > 29 + 1 5x > 30 x > 6

OpenStudy (thedj4jc):

That wouldn't work, but thanks for trying!!

OpenStudy (anonymous):

oh sorry not that good at math what grade are you in anyways?

OpenStudy (thedj4jc):

9th/10th, I'm doing Algebra 1

OpenStudy (anonymous):

can u help me with something

OpenStudy (thedj4jc):

I can try!!

OpenStudy (anonymous):

10 • [2 + (24 – 8)] ÷ 4

OpenStudy (thedj4jc):

Ok, so you're gonna start with the problem in the parentheses. 24-8=?

OpenStudy (anonymous):

16

OpenStudy (thedj4jc):

Ok, good! Now, 16+2=?

OpenStudy (anonymous):

18

OpenStudy (thedj4jc):

Yep!! Now, 10*180=?

OpenStudy (thedj4jc):

18*

OpenStudy (anonymous):

180

OpenStudy (thedj4jc):

Then, 180/4=?

OpenStudy (anonymous):

45 thank u so much

OpenStudy (thedj4jc):

Anytime!!!